OpenStudy (anonymous):

simplify the following fraction (please show work!).... (x/2x+3) - (4/5x-1)

OpenStudy (anonymous):

is it x 4 (--- + 3 ) - (---- - 1) ????? 2x 5x

OpenStudy (anonymous):

i have no idea what you just typed

OpenStudy (anonymous):

is this your question as printed in the book ???

OpenStudy (anonymous):

\[(x \div2x +3)-(4\div5x -1)\]

OpenStudy (anonymous):

ok ill ask again because it is still not clear is it x divide by (2x+3) minus 4 divided by (5x-1) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, then i know the right question...

OpenStudy (anonymous):

ok can you answer it plz

OpenStudy (anonymous):

x(5x-1) - 4(2x+3) = ------------------ (taking LCD) (2x+3) (5x-1) 5x^2 - x - 8x -12 = -------------------- 10x^2 - 2x + 15x - 3 5x^2 - 9x - 12 = --------------- 10x^2 + 13x - 3

OpenStudy (anonymous):

i cant understand this can you explain step by step

OpenStudy (anonymous):

since both terms r fractions, we have to take their LCD when denominators are algebraic terms which have no common factor visible, we take LCD as their product. Now we divide the LCD by the denominator of each fraction and multiply the numerator by the remainder.... (2x+3)(5x-1) ------------ leaves (5x-1) so we multiply the numerator x by (5x-1) (2x+3) (2x+3)(5x-1) ------------ leaves (2x+3) so we multiply the numerator 4 by (2x +3) (5x-1) then we resolve the brackets..... is it clear now????

OpenStudy (anonymous):

can you give me a final answer using the equation button

OpenStudy (anonymous):

\[5x ^{2} - 9x - 12\] -------------------\[10x ^{2} + 13x -3\]

OpenStudy (anonymous):

thanks i get what your saying now would you mind helping me with\[(2\div \cos \theta)+(4\div1-\sin \theta)\]

OpenStudy (anonymous):

let me see...... give a minute or two

OpenStudy (anonymous):

sorry, not able to find a proper way to do it, pls put it out as a new question for others....

OpenStudy (anonymous):

ok could you answer a diffrent question